OpenStudy (anonymous):

solution to the equation 25^(z – 4) = 125 ?

OpenStudy (anonymous):

You must use logarithms. danielcoolxx, 25 is being raised to a power, not multiplied; therefore your method is not correct.

OpenStudy (campbell_st):

re-write the equation in terms of powers of 5 \[(5^2)^{z - 4} = 5^3\] using the power of a power law \[5^{2x - 8} = 5^3\] then equate the powers 2x - 8 = 3 2x = 11 x = 11/2

OpenStudy (campbell_st):

oops changed variables from z to x... sorry about that